blems: Series A PR 1-1A Transactions OBJ. 4 1 1 of the current year, Andrea Byrd established a business to manage 85 property. She completed the following transactions during April: a. Opened a business bank account with a deposit of $45,000 from personal funds. b. Purchased office supplies on account, $2,000. c. Received cash from fees earned for managing rental property, $8,500 d. Paid rent on office and equipment for the month, $5,000. e. Paid creditors on account, $1,375 f. Billed customers for fees earned for managing rental property, $11,250. g. Paid automobile expenses (including rental charges) for month, $840, and miscella- neous expenses, $900. h. Paid office salaries, $3,600. i. Determined that the cost of supplies on hand was $550; therefore, the cost of supplies used was $1,450 j. Withdrew cash for personal use, $2,000. Instructions 1. Indicate the effect of each transaction and the balances after each transaction, using the following tabular headings
Chapter 1 l 1. Cash received from fees earned. 2. Cash paid for expenses 3. Cash paid for land. 4. Cash paid to owner for personal use.
Answer #1

Ans-

Assets = Liabilities + Owner's Equity
Item Cash + Accounts Receivables + Supplies = Accounts Payable + Andrea Byrd, Capital - Andrea Byrd, Drawings + Fees Earned - Rent Expenses - Salaries Expenses - Supplies Expenses - Auto Expenses - Miscellaneous Expenses
a. 45,000 + + = + 45,000 - + - - - - -
Balance 45,000 + + = + 45,000 - + - - - - -
b. + + 2,000 = 2,000 + - + - - - - -
Balance 45,000 + + 2,000 = 2,000 + 45,000 - + - - - - -
c. 8,500 + + = + - + 8,500 - - - - -
Balance 53,500 + + 2,000 = 2,000 + 45,000 - + 8,500 - - - - -
d. -5,000 + + = + - + - -5,000 - - - -
Balance 48,500 + + 2,000 = 2,000 + 45,000 - + 8,500 - -5,000 - - - -
e. -1,375 + + = -1,375 + - + - - - - -
Balance 47,125 + + 2,000 = 625 + 45,000 - + 8,500 - -5,000 - - - -
f. + 11,250 + = + - + 11,250 - - - - -
Balance 47,125 + 11,250 + 2,000 = 625 + 45,000 - + 19,750 - -5,000 - - - -
g. -1,740 + + = + - + - - - - -840 - -900
Balance 45,385 + 11,250 + 2,000 = 625 + 45,000 - + 19,750 - -5,000 - - - -840 - -900
h. -3,600 + + = + - + - - -3,600 - - -
Balance 41,785 + 11,250 + 2,000 = 625 + 45,000 - + 19,750 - -5,000 - -3,600 - - -840 - -900
i. + + -1,450 = + - + - - - -1,450 - -
Balance 41,785 + 11,250 + 550 = 625 + 45,000 - + 19,750 - -5,000 - -3,600 - -1,450 - -840 - -900
j. -2,000 + + = + - -2,000 + - - - - -
Balance 39,785 + 11,250 + 550 = 625 + 45,000 - -,2000 + 19,750 - -5,000 - -3,600 - -1,450 - -840 - -900